# Hybrid deep learning-based strategy for the hepatocellular carcinoma cancer grade classification of H&E stained liver histopathology images

### v1

@@ Abstract
Hepatocellular carcinoma (HCC) is a common type of liver cancer whose early-stage diagnosis is a common challenge, mainly due to the manual assessment of hematoxylin and eosin-stained whole slide images, which is a time-consuming process and may lead to variability in decision-making. For accurate detection of HCC, we propose a hybrid deep learning-based architecture that uses transfer learning to extract the features from pre-trained convolutional neural network (CNN) models and a classifier made up of a sequence of fully connected layers. This study uses a publicly available The Cancer Genome Atlas Hepatocellular Carcinoma (TCGA-LIHC)[1] database (n=491) for model development and proprietary database of Kasturba Gandhi Medical College (KMC), India [2] for validation. The pre-processing step involves patch extraction, colour normalization, and augmentation that results in 3920 patches for the TCGA dataset. The developed hybrid deep neural network consisting of a CNN-based pre-trained feature extractor and a customized artificial neural network-based classifier is trained using five-fold cross-validation. For this study, eight different state-of-the-art models are trained and tested as feature extractors for the proposed hybrid model.
The proposed hybrid model with ResNet50-based feature extractor provided the sensitivity, specificity, F1-score, accuracy, and AUC of 100.00%, 100.00%, 100.00%, 100.00%, and 1.00, respectively on the TCGA database. On the KMC database, EfficientNetb3 resulted in the optimal choice of the feature extractor giving sensitivity, specificity, F1-score, accuracy, and AUC of 96.97$\pm$0.83, 98.85$\pm$0.24, 96.71$\pm$0.66, 96.71$\pm$0.66, and 0.99$\pm$0.01, respectively. The proposed hybrid models showed improvement in accuracy of 2% and 4% over the pre-trained models in TCGA-LIHC and KMC databases.
@@ I Introduction
With the advancements in artificial intelligence (AI), the landscape of medical diagnosis has completely changed, and it has helped over the years the medical community in providing patients with accurate and inexpensive computer-aided diagnostic solutions for the betterment of their health [11]. In the domain of liver cancer detection, the AI-based convolutional neural network (CNN) algorithms have contributed to a great extent in the prediction of cancer types [6] [2][12][13][7] replacing traditional techniques such as SVM[14]. In the past decade, several CNN-based architectures have been proposed that either focus on deepening the network such as AlexNet[15] and VGG[16] or to overcome the vanishing gradient issue by adding the residual connections ResNet [17] architectures and DenseNet [18] or applying compound scaling methodology that demonstrates dependency on width, depth and resolution of CNN through EfficientNet architectures [19]. These models have been used as state-of-the-art for a long time and possess special features. Most of the recent studies used such CNN architecture as their backbone models. Unlike the conventional machine learning (ML)-based algorithms, the CNNs do not require handcrafted features for training. The architecture of CNN is responsible for both feature extraction and classification eventually achieving better performance than traditional methods[20][21] and providing an end-to-end solution. This is the primary reason for their popularity in performing medical image classification [22].
@@
Recently, Sun et al.[12] used a ResNet50to perform the liver cancer classification using global image-based labels. The authors proposed patch extraction and feature selection mechanisms. Chen et al.[13] used the InceptionV3 to perform the patch-level liver cancer classification on the TCGA dataset. Atresh et al.[2] proposed a new LiverNet architecture, which consisted of a convolutional block attention module block (CBAM), atrous convolution spatial pyramid pooling blocks, and hypercolumn technique, which reduces model parameters as well as makes the model more robust and precise. The authors claimed the overall liver cancer detection accuracy of 97.72%. In another study, Chen et al.[7] studied the performance of ResNet50, ResNet50_CBAM, SENet, and SKNet on purely, moderately and well-differentiated histopathology images. SENet and ResNet50 were found to be the best-performing models on poorly and well-differentiated types, respectively. Another study [23] proposes prior segmentation for better classification accuracies. Luo et al. [24] by employing a deformable convolution guided attention block and deep adaptive feature fusion, proposed DCA-DAFFNet for laryngeal tumor grading and achieves accuracy of 90.78%. Similarly, few more recent studies use the CNN-based algorithms for histopathology image classification [25][26] [20][27].
All these studies used the diverse CNN-based architecture for the liver cancer classification. Most of these studies also use two most important concepts i.e. transfer-learning and fine-tuning. In transfer learning, the model trained on the diversified and huge database with multiple classes is used to train on a small dataset with a limited number of classes[28][29] [30] [31]. Similarly, in fine-tuning, selective top layers of feature extractor along with the classifier are made trainable[25][32][33][26][34]. Recent study by Talo et al.[30] suggest that if some of the top layers of the feature extractor are kept trainable along with the classifier, then better classification accuracy can be achieved.
Motivated by the concept of transfer learning and fine-tuning, and observing that keeping selective top feature extractor layers along with classifier trainable can boost overall classification performance[25], we developed a hybrid CNN-based architecture for liver cancer classification. We hypothesized that the proposed hybrid deep learning-based architecture may perform better than the individual pre-trained models. The salient contributions are listed as follows:
@@
    Fine-tuning powered model is designed to address the issues in cross-domain learning. In the proposed work, the effect of transferring the learnings to the medical domain and tuning the model to a certain depth for new characterization is studied.
@@
    Classifier designed with fully connected layers with gradual feature reduction to ensure proper mapping to the output label space. This also helps to develop an optimum end-to-end solution for histopathological image classification.
@@
    This study also demonstrates the importance of proper image pre-processing and training methodology for better loss convergence, achieving optimum results and developing the robust model.
The rest of the paper is organized as follows. Section 2 discusses the details of the dataset, preprocessing, a proposed model architecture, performance metrics and training methodology. Section 3 describes the experimental details with results validation along with a detailed comparative study. In Section 4, a detailed discussion on the experimental results obtained, followed by the conclusion in Section 5. The overall workflow for the experimentation is shown in Fig. 1
@@ II-A Datasets and Pre-processing
This study uses a publicly available cancer genome atlas program liver hepatocellular carcinoma (TCGA-LIHC) database of 491 whole slide images [1] to train the proposed hybrid deep learning-based model. The 491 slides were divided into three types of liver HCC tumor namely solid tissue normal as type 0, primary tumor as type 1, and recurrent tumor as type 2. There were 89 solid tissue normal slides, 399 primary tumor slides, and 3 recurrent tumor slides in total. In a deep learning-based approach for classification, directly feeding the whole slide images of very large dimensions to CNN is nearly impractical as it increases computations and model parameters by a large amount. Downsampling or reshaping the histopathology images to lower dimensions leads to severe information loss. Therefore, we follow a patch-based approach [12], where each WSI slide is tiled into patches of size 1024$\times$1024. While performing the patch extraction, we also intended to maintain sufficient visibility of tissues in the patches. This was ensured by maintaining the mean intensity value of 200 and the standard deviation of 60. An empirical selection process is performed to maintain the clear visualization of tissues in the patches and to reduce the class imbalance in the database. Finally, 813, 893, and 680 valid patches of type 0, type 1 and type 2, respectively are considered to develop the proposed hybrid deep learning-based model.
@@
Figure 2: TCGA-LIHC liver HCC database[1]. This diagram describes the workflow overall TCGA dataset preparation which involves patch extraction from whole slide images, then downsampling to balance the dataset, and augmentation. After data preprocessing the number of images obtained are mentioned in the table.
@@
The color normalization is another important step in the pre-processing of histology images because the variation of staining intensity of hematoxylin and eosin used for staining purposes in the histology images by the pathologist and may vary according to the practitioner and physical conditions. Thus, the color variations can significantly impact the model training because models are colour-sensitive, In the proposed study, we used the color normalization technique proposed by Macenko et al.[35] which presents an algorithm that finds the correct stain vectors for the image after converting image pixels to optical density space automatically and then performs colour deconvolution.
@@
Deep learning models require a diverse and large dataset to build robust and flexible networks. Augmentation is a technique which generates variants of existing data ultimately increasing the dataset. This makes data augmentation an important step in preprocessing. Over the normalized patches, we applied random vertical and horizontal flips and obtained 1220, 1340, and 1360 patches of type 0, type 1, and type 2 respectively. The complete pre-processing stage is depicted in Fig. 2.
@@
This study also uses another cancer database of liver HCC collected from Kasturba Gandhi Medical College, India [2]. The KMC database had four types of liver HCC tumors such as type 0, type 1, type 2, and type 3 having 719, 799, 776, 711 image patches, respectively, each of size 224$\times$224$\times$3. The received KMC database[2], was originally pre-processed and, therefore, only the augmentation step is performed during the training of the proposed model. Moreover, to verify the proposed hybrid learning strategy over the different types of cancer disease, another publicly available colon cancer database of histopathology images is collected from Kaggle [36]. The colon database consists of two classes, adenocarcinoma and normal, with a total of 5000 patches per type of size 768$\times$768$\times$3. Samples of all three datasets are shown in Fig. 3
@@ II-B Proposed model architecture
Using transfer learning, we can apply models that have been trained on large diverse datasets (i.e., “pre-trained models”) to smaller, related datasets in a more efficient manner [25, 32, 33, 30, 29, 12]. In general, the pre-trained models have two parts: a feature extractor and a classifier. In transfer learning, the classifier is modified while the feature extractor is left untouched. This is because the pre-trained models are usually trained on the Imagenet dataset with 1000 output classes, while in general, the number of output classes is limited. Therefore, only the last classification layer of the pre-trained model is re-trained and named as “base model".
In general, the top layers of any pre-trained model’s feature extractor capture the fine details that pertain solely to output classes, and therefore, such top layers are re-trained when fine-tuning the model on small custom datasets to optimize the performance of the model [25, 32],[33]. In contrast to the top layers, the bottom layers capture the more generic features such as lines, edges, and shapes that contribute more to the classification task. Besides the feature extractor and the classifier, another important attribute of the pre-trained model is the depth of the model. There has been evidence that deeper models can improve classification performance, although this is not always the case [17, 18, 19]. Considering all these aspects, we propose an architecture in which (i) the feature extractor remains unchanged except for a few top layers and (ii) a few more fully connected layers are added before the final classification layer. This modified model is referred to as the “hybrid model”, which combines both the concepts of transfer learning and fine-tuning. The number of additional fully connected layers is decided empirically. The performance of eight different types of pre-trained models such as ResNet50[17], VGG16[16], EfficientNetb0[19], EfficientNetb1[19], EfficientNetb2[19], EfficientNetb3[19], EfficientNetb4[19], DenseNet121[18] is compared while designing the hybrid model. Fig. 4 shows the architecture of base and hybrid models.
Figure 4: i) base model: All of the convolutional layers in pre-trained model are frozen and the last fully connected layer is replaced and kept trainable. ii) hybrid model: Bottom layers of the convolutional of the pre-trained model are frozen. A shallow classifier is replaced with a deep classifier. classifier and selective top layers of the pre-trained model are kept trainable.
@@ II-C Training
The overall architecture of the proposed model is comprised of two stages: (i) offline system and (ii) online system as shown in Fig. 5. PyTorch framework was used for designing the model. Initially, the pre-processed image patches are divided into training (90%) and testing datasets (10%) as shown in Table I. The image patches are resized to 224$\times$224$\times$3 pixel dimensions to maintain uniformity across datasets and allow faster and more precise training. The offline system uses the training dataset and performs the 5-fold stratified cross-validation to train the proposed model. In 5-fold cross-validation, the training data is divided into five equal non-overlapping parts. The model is trained on the four parts and validated on the remaining part of the training dataset. Dynamic augmentation is performed on these four training parts of the data with random rotation on the TCGA-LIHC dataset, random rotation with random horizontal flip on the COLON dataset, and random rotation with random horizontal and vertical flip on the KMC dataset. Furthermore, to minimize the class imbalance, a weighted random sampler approach [37] is followed that samples images of different classes with pre-defined weights. The weights are inversely utilized based on the number of classes in the training dataset. Both the dynamic augmentation and weighted random sampler methods aid in reducing the class-imbalance effect in the dataset. This complete process is repeated five times in 5-fold cross-validation and an averaged validation performance is recorded. The hyperparameters used while training the model are shown in Table II. Moreover, The cosine annealing warm restart [38] learning rate scheduler is used to select the learning rate over all the epochs. This is mathematically represented in Eq.[1], where, $\eta^{i}_{min}$ and $\eta^{i}_{max}$ are learning rate ranges. $T_{cur}$ indicates the number of epochs from the last restart, and $T_{i}$ indicates the number of epochs after which a restart is scheduled. The model is trained for 47 epochs with early stopping criteria and a warm restart of the learning rate scheduler at every $12^{th}$ epoch($T_{i}$). In this study, an initial learning rate of 0.001($\eta^{i}_{max}$) is used.

### v2

# Hybrid deep learning-based strategy for hepatocellular carcinoma cancer grade classification

@@ Abstract
Hepatocellular carcinoma (HCC) is a common type of liver cancer whose early-stage diagnosis is a challenge. This is mainly done by a manual assessment of hematoxylin and eosin-stained whole slide images, which is a time-consuming process with high variability in the decision-making. In order to achieve accurate detection of HCC, we propose a hybrid deep learning architecture that leverages the advantages of fine-tuning selective layers of convolutional neural network (CNN) alongside a deep classifier comprising a sequence of fully connected layers, i.e., Artificial Neural Network (ANN). In order to validate the method, this study uses a publicly available The Cancer Genome Atlas Hepatocellular Carcinoma (TCGA-LIHC)[1] database (n=491) for model development and proprietary database of Kasturba Gandhi Medical College (KMC), India [2] for validation. The proposed hybrid model with ResNet50-based feature extractor provided the sensitivity, specificity, F1-score, accuracy, and AUC of 100.00%, 100.00%, 100.00%, 100.00%, and 1.00, respectively on the TCGA database. It was contrasted with eight different state-of-the-art models. On the KMC database, EfficientNetb3 resulted in the optimal choice of the feature extractor giving sensitivity, specificity, F1-score, accuracy, and AUC of 96.97$\pm$0.83, 98.85$\pm$0.24, 96.71$\pm$0.66, 96.71$\pm$0.66, and 0.99$\pm$0.01, respectively. The proposed hybrid models showed improvement in accuracy of 2% and 4% over the pre-trained models in TCGA-LIHC and KMC databases.
@@ I Introduction
With advances in artificial intelligence (AI), the landscape of medical diagnosis has completely changed, and it has helped over the years the medical community in providing patients with accurate and inexpensive computer-aided diagnostic solutions for the betterment of their health [11]. In the domain of liver cancer detection, the AI-based convolutional neural network (CNN) algorithms have contributed to a great extent to the prediction of cancer types [6] [2][12][13][7] replacing traditional techniques such as SVM[14]. In the past decade, several CNN-based architectures have been proposed that either focus on deepening the network such as AlexNet[15] and VGG[16], overcoming the vanishing gradient issue by adding the residual connections, ResNet [17] architectures and DenseNet [18] or applying compound scaling methodology that demonstrates dependency on width, depth and resolution of CNN through EfficientNet architectures [19]. These models have been used as state-of-the-art for a long time and possess special features. Most of the recent studies used this CNN architecture as their backbone models. Unlike the conventional machine learning (ML)–based algorithms, the CNNs do not require handcrafted features for training. The architecture of CNN is responsible for both feature extraction and classification, eventually achieving better performance than traditional methods[20][21] and providing an end-to-end solution. This is main reason for their popularity in performing medical image classification [22].
@@
Recently, Sun et al.[12] used a ResNet50 to perform the liver cancer classification using global image-based labels. The authors proposed patch extraction and feature selection mechanisms. Chen et al.[13] used the InceptionV3 to perform the patch-level liver cancer classification on the TCGA dataset. Atresh et al.[2] proposed a new LiverNet architecture, which consisted of a convolutional block attention module block (CBAM), atrous convolution spatial pyramid pooling blocks, and hypercolumn technique, which reduces model parameters as well as makes the model more robust and precise. The authors claimed an overall liver cancer detection accuracy of 97.72%. In another study, Chen et al.[7] studied the performance of ResNet50, ResNet50_CBAM, SENet, and SKNet on purely, moderately and well-differentiated histopathology images. SENet and ResNet50 were found to be the best-performing models on poorly and well-differentiated types, respectively. Another study [23] proposes prior segmentation for better classification accuracy. Luo et al. [24] employed a deformable convolution-guided attention block and deep adaptive feature fusion, and proposed DCA-DAFFNet for laryngeal tumor grading achieving an accuracy of 90.78%. Similarly, few more recent studies used the CNN-based algorithms for histopathology image classification [25][26] [20][27].
All these studies used diverse CNN-based architecture for liver cancer classification. Most of these studies relied on transfer learning, particularly fine-tuning, which is its most widely used form. In transfer learning, the model trained on a diversified and huge database with multiple classes is used to train on a smaller dataset with a limited number of classes[28][29] [30] [31]. Similarly, in fine-tuning, selective top layers of feature extractor together with the classifier are made trainable[25][32][33][26][34]. A recent study by Talo et al.[30] suggested that if some of the top layers of the feature extractor are kept trainable along with the classifier, then better classification accuracy can be achieved.
Motivated by the concept of transfer learning, and observing that keeping selective top feature extractor layers along with classifier trainable can boost overall classification performance[25], we developed a hybrid CNN-based architecture for liver cancer classification. We hypothesized that the proposed hybrid deep learning-based architecture performs better than the individual pre-trained models. Hence, our contributions here are as follows:
@@
    We propose a deep classifier designed with gradual feature reduction to ensure proper mapping to the output label space. This also helps to develop an optimum end-to-end solution for histopathological image classification.
@@
    We applied a fine-tuning-powered model to address the issues in cross-domain learning. In the proposed work, the effect of transferring the learning to the medical domain and selectively tuning the model for new characterization is studied.
@@
    This study also demonstrates the importance of proper image pre-processing and training methodology for better loss convergence, achieving optimal results and developing a robust model for carcinoma detection in histopathological images.
The rest of the paper is organized as follows. Section 2 discusses the details of the dataset, the preprocessing, the proposed model architecture, the performance metrics and the training methodology. Section 3 describes the experimental details within the validation along with a detailed comparative study. In Section 4, a detailed discussion is given of the experimental results obtained, followed by the conclusions in Section 5. The overall workflow for the experimentation is shown in Fig. 1.
@@ II-A Datasets and Pre-processing
This study uses a publicly available cancer genome atlas program liver hepatocellular carcinoma (TCGA-LIHC) database of 491 whole slide images [1] to train our proposed hybrid deep learning-based model. The 491 slides were divided into three types of liver HCC tumor namely solid tissue normal as type 0, primary tumor as type 1, and recurrent tumor as type 2. There were 89 solid tissue normal slides, 399 primary tumor slides, and 3 recurrent tumor slides. In a deep learning-based approach for classification, directly feeding the whole slide images of huge dimensions to CNN is nearly impractical as it increases computations and model parameters by a large amount. Downsampling or reshaping the histopathology images to lower dimensions leads to severe information loss. Therefore, we follow a patch-based approach [12], where each WSI slide is tiled into patches of size 1024$\times$1024. While performing the patch extraction, we also intended to maintain sufficient visibility of tissues in the patches. This was ensured by maintaining the mean intensity value of 200 and the standard deviation of 60. An empirical selection process is performed to maintain a clear visualization of tissues in the patches and reduce the class imbalance in the database. Finally, this process produced 813, 893, and 680 valid patches of type 0, type 1, and type 2, respectively, which are considered for developing the hybrid deep learning-based model.
@@
Figure 2: TCGA-LIHC liver HCC database[1]. This diagram describes the preprocessing for TCGA dataset preparation, which involves patch extraction from the whole slide image, downsampling to balance the dataset, and data augmentation. After data preprocessing, the number of images obtained are mentioned at the bottom.
@@
The color normalization is another important step in the pre-processing of histology images because of the variation of staining intensity of hematoxylin and eosin used for staining purposes in the histology images by the pathologist and may vary according to the practitioner and physical conditions. Thus, color variations can significantly impact model training, because models are color-sensitive, In the proposed study, we used the color normalization technique proposed by Macenko et al.[35] which presents an algorithm that finds the correct stain vectors for the image after converting image pixels to optical density space automatically and then performs color deconvolution.
@@
Deep learning models require a diverse and large dataset to build robust and flexible networks. Augmentation is a technique that generates variants of existing data, ultimately increasing the dataset. This makes data augmentation an important step in preprocessing. Over the normalized patches, we applied random vertical and horizontal flips and obtained 1220, 1340, and 1360 patches of type 0, type 1, and type 2 respectively. The complete pre-processing stage is depicted in Fig. 2.
@@
This study also uses another cancer database of liver HCC collected from Kasturba Gandhi Medical College, India [2]. The KMC database has four types of liver HCC tumors such as type 0, type 1, type 2, and type 3 having 719, 799, 776, and 711 image patches, respectively, each of size 224$\times$224$\times$3. The KMC database received[2], was originally pre-processed and, therefore, only the augmentation step is performed during the training of the proposed model. Moreover, to verify the proposed hybrid learning strategy for different types of cancer disease, another publicly available colon cancer database of histopathology images is collected from Kaggle [36]. The colon database consists of two classes, adenocarcinoma and normal, with a total of 5000 patches per type of size 768$\times$768$\times$3. Samples of all three datasets are shown in Fig. 3
@@ II-B Proposed model architecture
The transfer learning property of NNs allows to apply models that have been trained on large diverse datasets (i.e., “pre-trained models”) to smaller datasets from completely different domains in a more efficient manner [25, 32, 33, 30, 29, 12]. In general, the pre-trained models have two parts: a feature extractor and a classifier. In transfer learning, the classifier is modified while the feature extractor is left untouched. This is because the pre-trained models are usually trained on the Imagenet dataset with 1000 output classes, while in general, the number of output classes is limited. Therefore, only the last classification layer of the pre-trained model is re-trained, named as “base model".
In general, the top layers of any pre-trained model’s feature extractor capture the fine details that pertain solely to output classes, and therefore, such top layers are re-trained when fine-tuning the model on small custom datasets to optimize the performance of the model [25, 32],[33]. In contrast to the top layers, the bottom layers capture the more generic features such as lines, edges, and shapes that contribute more to the classification task. Besides the feature extractor and the classifier, another important attribute of the pre-trained model is the depth of the model. There has been evidence that deeper models can improve classification performance, although this is not always the case [17, 18, 19]. Considering all these aspects, we propose an architecture in which (i) the feature extractor remains unchanged except for a few top layers and (ii) a few more fully connected layers (ANN) are added before the final classification layer. This modified model, referred to as the "hybrid model", leverages the advantages of selective fine-tuning and deep classifiers. The number of additional fully connected layers is decided empirically. The performance of eight different types of pre-trained models such as ResNet50[17], VGG16[16], EfficientNetb0[19], EfficientNetb1[19], EfficientNetb2[19], EfficientNetb3[19], EfficientNetb4[19], DenseNet121[18] is compared while designing the hybrid model. Fig. 4 shows the architecture of the base and hybrid models.
Figure 4: i) Base model: All of the convolutional layers in pre-trained model are frozen and the last fully connected layer is replaced and kept trainable. ii) Hybrid model: Bottom layers of the convolutional layers of the pre-trained model are frozen. A shallow classifier is replaced with a deep classifier. The classifier and selective top layers of the pre-trained model are kept trainable.
The overall architecture of the proposed model is comprised of two stages: (i) offline system and (ii) online system as shown in Fig. 5.
@@ II-C Training
The image patches are resized to 224$\times$224$\times$3 pixel dimensions to maintain uniformity across datasets and allow faster and more precise training. The offline system uses the training dataset and performs the 5-fold stratified cross-validation to train the proposed model. In 5-fold cross-validation, the training data is divided into five equal non-overlapping parts. The model is trained on the four parts and validated on the remaining part of the training dataset. Dynamic augmentation is performed on these four training parts of the data with random rotation on the TCGA-LIHC dataset, random rotation with random horizontal flip on the COLON dataset, and random rotation with random horizontal and vertical flip on the KMC dataset. Furthermore, to minimize the class imbalance, a weighted random sampler approach [37] is followed that samples images of different classes with pre-defined weights. The weights are inversely utilized based on the number of classes in the training dataset. Both the dynamic augmentation and weighted random sampler methods aid in reducing the class-imbalance effect in the dataset. This complete process is repeated five times in 5-fold cross-validation and an averaged validation performance is recorded. Table II shows the hyperparameters used while training the model.
Furthermore, cross-entropy loss $(L_{CE})$ between labels and ground truth values is used, which is defined in Eq.[1]
$L_{CE}=-\sum_{i=0}^{N-1}y_{i}\log(p_{i})$ (1)
where $N$ is the number of classes, and $y_{i}$ and $p_{i}$ are the ground truth values and the predicted probabilistic values, respectively. This loss function is widely preferred over the others. The reason for this is that, if we see the curve of loss vs. probability, loss is very high when prediction is inaccurate. As a result, the gradient of loss is high, which results in faster convergence. Once the model is trained, the training weights are used to transform the test dataset into the output cancer types using the test model. The performance of the model is evaluated on the test data using five performance measures such as accuracy, sensitivity, specificity, F1-score, and area under the curve (AUC). The pre-trained models were imported from the Pytorch library and trained on Google Colab.
III Validation
In this section, we present the model training settings, performance metrics, and results obtained on the three datasets, followed by detailed discussions.
III-A Settings
PyTorch framework was used for designing the model. Initially, the pre-processed image patches are divided into training (90%) and testing datasets (10%) as shown in Table I. Cosine annealing warm restart [38] learning rate scheduler is used to select the learning rate over all epochs. This is mathematically represented in Eq.[2], where, $\eta^{i}_{min}$ and $\eta^{i}_{max}$ are learning rate ranges. $T_{cur}$ indicates the number of epochs from the last restart, and $T_{i}$ indicates the number of epochs after which restart is scheduled. The model is trained for 47 epochs with early stopping criteria and a warm restart of the learning rate scheduler at every $12^{th}$ epoch($T_{i}$). In this study, an initial learning rate of 0.001($\eta^{i}_{max}$) is used: